3(x-6)+6=5(x+2)-12
Answer by mmm4444bot(95)

Hello There:
3*(x - 6) + 6 = 5*(x + 2) - 12
(The asterisks mean multiplication.)
Start by using the distributive property to get rid of the parentheses.
3*(x - 6) means that both the x and the -6 inside of the parentheses gets multiplied by 3. Likewise, both the x and the 2 get multiplied by 5 on the other set of parentheses.
3*x - 18 + 6 = 5*x + 10 - 12
Do the arithmetic on each side to add the constants together.
3*x - 12 = 5*x - 2
Next, we want all of the x-terms to be on one side of the equals sign and the constants on the other side.
Let's subtract 3*x from both sides.
3*x - 3*x - 12 = 5*x - 3*x - 2
-12 = 2*x - 2
Add 2 to each side.
2 - 12 = 2*x = 2 - 2
-10 = 2*x
Divide both sides by 2 to solve for x.
-10/2 = (2/2)*x
x = -5
Let's check our result. We will substitute -5 for x into the original equation and simplify by doing the arithmetic to see if we end up with a true statement.
3*(x - 6) + 6 = 5*(x + 2) - 12
3*(-5 - 6) + 6 = 5*(-5 + 2) - 12
3*(-11) + 6 = 5*(-3) - 12
-33 + 6 = -15 - 12
-27 = -27
This is a true statement, so our solution is correct.
~ Mark